Interest rate. In 1972, Bob purchased a new Datsun 240Z for $2,600. Datsun later changed its name to Nissan, and the 1972 Datsun 240Z became a classic. Bolb kept his car in excellent condition and in 2002 could sell the carfor six times what he originally paid. What was Bobs annualized rate of return for the 30 years he owned this car? If he keeps the car for another thirty years and earns the same rate, what could he sell the car for in 2032? What was Bobs annualized rate of return for the 30 years he owned this car? % (Round to two decimal places.)
Answer #1

Bob's Initial Purchase (by the year 1972) cost of the Car = $2,600

Value/Price of the Car after 30 Years = $2,600 x 6 = $15,600

Annualized Rate of Return (ARR) = $(15,600/2,600)/30*100 = 20%

Value of the Car after another 30 years period would be;

ARR = (FV / PV)1 / Y  - 1

=> 20/100 = (FV/15600)1/30 - 1

= $3,703,070.50